OpenStudy (anonymous):

Find the six trigonometric functions of each of the following radians. (a) θ=3π/2 (b) θ=5π/4 (c) θ=−π/3 Find the six trigonometric functions of each of the following radians. (a) θ=3π/2 (b) θ=5π/4 (c) θ=−π/3 @Mathematics

OpenStudy (anonymous):

yeah

myininaya (myininaya):

\[\sin(\frac{3\pi}{2})=-1\]

myininaya (myininaya):

all you have to do is use the unit circle

OpenStudy (anonymous):

the six trigonometric functions we are to find is the sine,cosine,tagent,cosec,sec and cotangent

myininaya (myininaya):

\[\cos(\frac{3\pi}{2})=0\]

myininaya (myininaya):

\[\tan(\frac{3\pi}{2})=\frac{\sin(\frac{3\pi}{2})}{\cos(\frac{3\pi}{2})}=\frac{-1}{0} DNE\]

myininaya (myininaya):

\[\cot(\frac{3\pi}{2})=\frac{\cos(\frac{3\pi}{2})}{\sin(\frac{3\pi}{2})}=\frac{0}{-1}=0\]

myininaya (myininaya):

\[\csc(\frac{3\pi}{2})=\frac{1}{\sin(\frac{3\pi}{2})}=\frac{1}{-1}=-1\]

myininaya (myininaya):

\[\sec(\frac{3\pi}{2})=\frac{1}{\cos(\frac{3\pi}{2})}=\frac{1}{0} DNE\]

OpenStudy (anonymous):

i thought u will 1st find the cordinates that correspondes to each one on a unit circle,and then solve with it?

myininaya (myininaya):

yep you need to find sine and cosine like i did above

myininaya (myininaya):

all the other trig functions can either be written in terms of sine or cosine or sine and cosine

myininaya (myininaya):

sine and cosine are given to you on the unit circle

OpenStudy (anonymous):

pls teach me,how is it determined on the unit circle whether is root2/2 or -1 0r 0 is confusing to me

myininaya (myininaya):

the first coordinate is the cosine of the angle the second coordinate is the sine of the angle

OpenStudy (anonymous):

ok the important thing is finding the cosine and sine,tthen i can derive the rest using cosine and sine as their inverse?

myininaya (myininaya):

inverse?

OpenStudy (anonymous):

i mean reciprocal?

myininaya (myininaya):

you can find csc(theta) by doing the reciprocal of sin(theta) and toy can find sec(theta) by doing reciprocal of cos(theta)

OpenStudy (anonymous):

yeah sorry i meant reciprocal,thanks

myininaya (myininaya):

\[\tan(\theta)=\frac{\sin(\theta)}{\cos(\theta)}\]

myininaya (myininaya):

\[\cot(\theta)=\frac{\cos(\theta)}{\sin(\theta)}\]

OpenStudy (anonymous):

whats sine and cosine of 5pie/4 in unit circle?